OpenStudy (anonymous):

15. Solve: 3(x - 2) = -27 A. x = -9 B. x = -7 C. x = 7 D. x = -11 16. Solve: 7x - 4 - 2x + 11 = 47 A. x = 8 B. x = 6.9 C. x = 10.8 D. x = -8 17. Solve: -4(1 - 3x) = 20 A. x = 2 B. x = -8 C. x = -2 D. x = 0

OpenStudy (anonymous):

15. multiply in the 3, simplify then solve for x 3x-6 = -27 3x= -21 x=-7 16 simplify and solve for x 5x+7 =47 5x = 40 x=8 17.factor in the -4, simplify and solve for x -4+12x = 20 12x = 24 x = 2